Question 10:

Please estimate the cost of common equity.

You have been provided with the following data: D0 = $0.80; P0 = $22.50; and gL = 8.00% (constant). Based on the dividend growth model, what is the cost of common from reinvested earnings?
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = Div0 * (1 + g) = $0.80 * (1 + 8%) = $0.864

22.50 = 0.864 r-008

r - 0.08 = 0.0384

Cost of Equity, r = 11.84%